# This draft: December 17, 2025 First draft: March 5, 2025

### v1

@@ C.4.3 Portfolio turnover and real-world implementability
This structure follows standard execution-cost decompositions emphasizing effective bid–ask spreads and market impact as primary sources of trading frictions (e.g., bessembinder2003trade; frazzini2012trading). We evaluate representative hyperparameter values $\lambda\in\{0,0.3,0.5,0.7,1.0\}$ under four cost scenarios (0, 25, 50, 75 bps). Table C.5 summarizes the resulting performance measures.
@@
Table C.5: Robustness of CB-APM Long–Short Portfolio Performance to Transaction Costs.
This table reports portfolio performance for representative hyperparameter values ($\lambda\in\{0,0.3,0.5,0.7,1.0\}$) under four proportional transaction-cost assumptions: 0 bps, 25 bps, 50 bps, and 75 bps.
@@ C.5 Ablation studies
C.5.1 Effect of macroeconomic feature compression
Figure C.8: Out-of-Sample $R^{2}$ without Macroeconomic State Embeddings.
@@ C.5.1 Effect of macroeconomic feature compression
To evaluate the contribution of macroeconomic state embeddings to CB-APM’s performance, we conduct an ablation study by re-estimating the model without the autoencoder component. Figure C.8 reports the out-of-sample $R^{2}$ for annual stock return prediction (left) and consensus variable approximation (right) across varying values of the hyperparameter $\lambda$.
@@
Importantly, consensus approximation remains largely unaffected in this ablated model, as shown in the right panel of Figure C.8. While the model continues to reconstruct analysts’ consensus variables with reasonable accuracy, the absence of macroeconomic embeddings severs an important informational channel linking consensus to return-relevant fundamentals. This divergence underscores the complementary function of the autoencoder that by distilling high-dimensional macroeconomic signals into latent state variables, it enriches the consensus layer with persistent economy-wide information, thereby mitigating noise in firm-level predictors and improving the model’s capacity to extract long-horizon risk premiums.
C.5.2 Role of joint optimization in consensus learning
@@
Table C.6 reports annual out-of-sample $R^{2}$ for this consensus-only specification, separately for a range of consensus-based targets and for returns. Consensus-related $R^{2}$ measures such as EPS forecast revision, Earnings forecast revisions, and Analyst Value remain economically sizable and relatively stable over time, while Analyst earnings per share stays in a narrow band around $80-86\%$ and EPS Forecast Dispersion between roughly $41\%$ and $52\%$. The composite Consensus average fluctuates only modestly between $27.52\%$ and $31.72\%$ across 2014–2023, with a full-sample value of $30.30\%$, indicating that the model recovers a stable consensus structure even without any return signal.
@@
Table C.6: Annual out-of-sample $R^{2}$ (%) under consensus-only training ($\lambda\to\infty$).
@@
Appendix D Detailed Data Description
D.1 Firm-level predictors
Table D.1: Descriptions of firm-level predictors from ChenZimmermann2021.
@@ 99 DebtIssuance Debt Issuance Spiess and Affleck-Graves 1999 JFE Annual
Table D.1: Descriptions of firm-level predictors from ChenZimmermann2021 (cont’d).
D.2 Macroeconomic predictors
Table D.2: Descriptions of macroeconomic predictors from welch2008comprehensive.
@@ D.2 Macroeconomic predictors
Table D.2: Descriptions of macroeconomic predictors from welch2008comprehensive (Cont’d).
Table D.3: Descriptions of macroeconomic predictors from FRED-MD (mccracken2016fred).
@@ 99 Prices CPIAUCSL CPI : All Items
Table D.3: Descriptions of macroeconomic predictors from FRED-MD (mccracken2016fred) (cont’d).
D.3 Analysts’ consensus variables
Table D.4: Descriptions of analysts’ consensus variables from ChenZimmermann2021.
@@ 9 AOP Analyst Optimism Frankel and Lee 1998 JAE
Table D.4: Descriptions of analysts’ consensus variables from ChenZimmermann2021 (cont’d).

### v2

@@ D.4.3 Portfolio turnover and real-world implementability
This structure follows standard execution-cost decompositions emphasizing effective bid–ask spreads and market impact as primary sources of trading frictions (e.g., bessembinder2003trade; frazzini2012trading). We evaluate representative hyperparameter values $\lambda\in\{0,0.3,0.5,0.7,1.0\}$ under four cost scenarios (0, 25, 50, 75 bps). Table D.5 summarizes the resulting performance measures.
@@
Table D.5: Robustness of CB-APM Long–Short Portfolio Performance to Transaction Costs.
This table reports portfolio performance for representative hyperparameter values ($\lambda\in\{0,0.3,0.5,0.7,1.0\}$) under four proportional transaction-cost assumptions: 0, 25, 50, and 75 bps.
@@ D.5 Ablation studies
D.5.1 Effect of macroeconomic feature compression
Figure D.5: Out-of-Sample $R^{2}$ without Macroeconomic State Embeddings.
@@ D.5.1 Effect of macroeconomic feature compression
To evaluate the contribution of macroeconomic state embeddings to CB-APM’s performance, we conduct an ablation study by re-estimating the model without the autoencoder component. Figure D.5 reports the out-of-sample $R^{2}$ for annual stock return prediction (left) and consensus variable approximation (right) across varying values of the hyperparameter $\lambda$.
@@
Importantly, consensus approximation remains largely unaffected in this ablated model, as shown in the right panel of Figure D.5. While the model continues to reconstruct analysts’ consensus variables with reasonable accuracy, the absence of macroeconomic embeddings severs an important informational channel linking consensus to return-relevant fundamentals. This divergence underscores the complementary function of the autoencoder that by distilling high-dimensional macroeconomic signals into latent state variables, it enriches the consensus layer with persistent economy-wide information, thereby mitigating noise in firm-level predictors and improving the model’s capacity to extract long-horizon risk premiums.
D.5.2 Avoiding look-ahead of AI analysts
@@
Table D.6 reports annual out-of-sample $R^{2}$ for this consensus-only specification, separately for a range of consensus-based targets and for returns. Consensus-related $R^{2}$ measures such as EPS forecast revision, Earnings forecast revisions, and Analyst Value remain economically sizable and relatively stable over time, while Analyst earnings per share stays in a narrow band around $80-86\%$ and EPS Forecast Dispersion between roughly $41\%$ and $52\%$. The composite Consensus average fluctuates only modestly between $27.52\%$ and $31.72\%$ across 2014–2023, with a full-sample value of $30.30\%$, indicating that the model recovers a stable consensus structure even without any return signal.
@@
Table D.6: Annual out-of-sample $R^{2}$ (%) under consensus-only training ($\lambda\to\infty$).
@@
Appendix E Detailed Data Description
E.1 Firm-level predictors
Table E.1: Descriptions of firm-level predictors from ChenZimmermann2021.
@@ 99 DebtIssuance Debt Issuance Spiess and Affleck-Graves 1999 JFE Annual
Table E.1: Descriptions of firm-level predictors from ChenZimmermann2021 (cont’d).
E.2 Macroeconomic predictors
Table E.2: Descriptions of macroeconomic predictors from welch2008comprehensive.
@@ E.2 Macroeconomic predictors
Table E.2: Descriptions of macroeconomic predictors from welch2008comprehensive (Cont’d).
Table E.3: Descriptions of macroeconomic predictors from FRED-MD (mccracken2016fred).
@@ 99 Prices CPIAUCSL CPI : All Items
Table E.3: Descriptions of macroeconomic predictors from FRED-MD (mccracken2016fred) (cont’d).
E.3 Analysts’ consensus variables
Table E.4: Descriptions of analysts’ consensus variables from ChenZimmermann2021.
@@ 9 AOP Analyst Optimism Frankel and Lee 1998 JAE
Table E.4: Descriptions of analysts’ consensus variables from ChenZimmermann2021 (cont’d).